# JigsawComm: Joint Semantic Feature Encoding and Transmission for Communication-Efficient Cooperative Perception

### v1

@@ 4.3 Ablation Study
Components OPV2V DAIR-V2X
Semantic
Loss
FUE &
Sched.
mAP@0.5 Size (KB) mAP@0.5 Size (KB)
- 0.93 26,292.36 0.55 22,670.62
- 0.94 13,516.80 0.61 7,374.995
0.92 43.84 0.69 279.97
Impact of Core Components. Table 3 ablates our two main contributions: the Semantic Loss $\mathcal{L}_{\text{semantic}}$ and the FUE & Scheduling module. Removing the FUE & Scheduler still allows the encoder’s $\mathcal{L}_{\text{semantic}}$ to create sparse feature maps. However, the entire maps are transmitted, leading to redundancy and a large payload. Without $\mathcal{L}_{\text{semantic}}$, the FUE & Scheduler deal with dense and semantically-overloaded features, making it hard to select due to noisy utility estimates, still resulting in a large payload. Only with the presence of both components, the transmission size is greatly reduced while attaining high CP accuracy.
Effect of Redundancy. The system’s transmission model is based on Theorem (1), demonstrating that selecting features from more than the Top-1 agent per cell is suboptimal. Fig. 8 validates this empirically. Comparing our Top-1 Agent policy to a Top-2 Agents policy, results show that Top-2 is worse at nearly every bandwidth point, indicating that transmitting redundant features wastes bandwidth and degrades perception accuracy. Although adding more data seems beneficial, this highlights a limitation in current fusion networks. Sec. 5 discusses this phenomenon and its implications further.
@@ 5 Discussion
On the ‘Cost’ of Redundancy. A critical, counterintuitive finding from our evaluation is that transmitting from the Top-2 agents per cell yields worse performance than transmitting the Top-1 agent alone, as shown in Fig. 8. This ‘more is worse’ outcome stems from a fundamental limitation presented in current SOTA CP fusion algorithms, such as element-wise max-pooling or dot-product attention [xu2022opv2v, hu2022where2comm, f_cooper]. The fused feature is a weighted sum of the received features, with the informative activations in the fused feature upper-bounded by the maximum signal across the individual features. Therefore, unlike human vision, these methods struggle to synergetically fuse redundant information. If the second-best agent’s feature is noisier or semantically different (e.g. conflicting channel activations), it can ‘dilute’ the clean, high-utility representation from the Top-1 agent. This finding corroborates the design philosophy of JigsawComm, which selects only the highest-quality, non-redundant data, resulting in smaller overhead and higher accuracy.
(a) OPV2V
(b) DAIR-V2X
Figure 8: Ablation on number of agents sharing per cell. Including the agent with the second-highest utility does not improve the utility-bandwidth tradeoff, which corroborates Theorem (1).
Why Not a Naive Communication Loss? A straightforward approach to reduce data size might be to add a simple communication loss directly on the transmission masks $\{\mathbf{M}\}_{i=1}^{N}$, penalizing the total number of transmitted bits. We found this approach to produce suboptimal performance in practice. Such a loss disregards the semantic importance and redundancy of the information transmitted, yet penalizes all communication. Our framework avoids this by training the FUE head to learn a task-aligned utility proxy, ensuring that selection decisions are driven by semantic-importance and redundancy.
Relation to Semantic Communications. JigsawComm is related to semantic/goal-oriented communications [strinati2021_6g, binucci2022adaptive, pezone2022goal], since they share the high-level goal of task-oriented data compression. However, existing work in semantic communications targets different applications [strinati2021_6g], considers only single sensor perception [yuan2023scalable, binucci2022adaptive, pezone2022goal], or neglects sensor correlations [sheng2024semantic]. Applying semantic communications to CP involves unique challenges due to its collaborative and multi-source input nature, where inter-agent redundancy prevails due to overlapping sensory data. Essentially, the CP problem studied in this paper can be reformulated as a joint, multi-source, task-oriented data compression problem, which is difficult to solve without knowing a priori the exact inter-agent data correlation structure. JigsawComm solves this problem via a data-driven approach (with a small overhead of exchanging semantic feature utilities).
Meanwhile, the semantic regularization and redundancy-eliminating transmission scheduling design of JigsawComm can be viewed as a way of implementing the information bottleneck (IB) principle, where unnecessary information is minimized while task-relevant information is maximized, which is essential to achieve goal-oriented communications [strinati2021_6g, pezone2022goal]. Interestingly, the same IB principle has also been shown to be effective in reducing generalization errors in deep learning models [kawaguchi2023does]. Here, our end-to-end trained JigsawComm learns to keep only semantically-relevant information from its sparse encoder and selects non-redundant and high-utility features for transmission, which effectively controls the IB of the CP pipeline. Therefore, in addition to the drastic reduction in required bandwidth for real-time communication, JigsawComm also achieves matching or improved perceptual quality compared to full-feature baselines.
Practical Implementation. Our framework is designed for practical V2X implementation by leveraging its dual-channel architecture [FCC_DSRC, CV2X_ETSI_EN_303_613]. The meta utility maps, though scaling linearly with agent count, can be transmitted over the control channel, incurring negligible overhead (<<0.03 KB per agent under 4-bit quantization). The final, sparse BEV features are then sent efficiently over the data channel, such as via the Collective Perception Message (CPM) defined by ITS [cpm]. Note that, our scheduler outputs the transmission priorities of all agents’ selected features as a byproduct (ordered by the corresponding feature utility values), which can be easily integrated into the low-layer V2X protocols for both scheduling-based (e.g. C-V2X, 5G NR-V2X [CV2X_ETSI_EN_303_613, garcia2021tutorial]) and random-access V2X protocols (e.g. DSRC [FCC_DSRC]). Our scheme is compatible with both centralized (V2I) and decentralized (V2V) operating modes. For V2I, an edge server co-located with the RSU can perform scheduling, fusion, and inference. For V2V, the same set of computations is performed by each agent in a distributed manner.
@@ 6 Conclusion and Future Work
We present JigsawComm, a novel joint semantic feature encoding and transmission framework for CP that maximizes every transmitted bit’s contribution to the final task accuracy. By jointly training a semantic sparse encoder and a lightweight Feature Utility Estimator via a differentiable scheduler, JigsawComm learns to extract and select only essential, non-redundant data. This enables a provably optimal top-1 policy at inference that eliminates cross-agent redundancy and achieves a scalable $\mathcal{O}(1)$ communication cost. Our experiments on OPV2V and DAIR-V2X demonstrate SOTA accuracy with up to over $500\times$ bandwidth reduction.
There are several promising directions for future work. First, as JigsawComm adopts the index-based encoding to convert sparsely encoded features to bits, we will further push the limit of JigsawComm’s bandwidth efficiency by integrating traditional lossless source coding methods such as entropy coding [storer1988data]. Second, we will also extend JigsawComm with joint semantic source and channel coding to deal with unreliable channel conditions. Third, we will study the impact of different sensor modalities and feature cell granularities on the accuracy-bandwidth tradeoff, and explore novel fusion networks that can leverage, rather than suffer from, redundancy. Finally, we will study practical V2X-based CP system design, robustness to communication failures or packet losses, and end-to-end goal-oriented CP systems for autonomous driving.

### v2

@@ 4.4 Ablation Study
OPV2V DAIR-V2X
$\mathcal{L}_{\text{sem}}$ FUE mAP@0.5 Size (KB) Rel. Eff. (%) mAP@0.5 Size (KB) Rel. Eff. (%)
0.93 26,292.4 0.17 0.55 22,670.6 0.99
0.94 13,516.8 0.33 0.61 7,375.0 3.38
0.92 43.8 100 0.69 282.0 100
Impact of Core Components. Table 2 ablates $\mathcal{L}_{\text{semantic}}$ and the FUE & Scheduler. Using $\mathcal{L}_{\text{semantic}}$ alone produces sparse features but transmits them in full, yielding 26,292 KB on OPV2V because inter-agent redundancy is unaddressed. Using the FUE alone selects features intelligently but operates on dense representations (13,517 KB). Only when both are combined does transmission drop to 44 KB while maintaining high accuracy: $\mathcal{L}_{\text{semantic}}$ ensures each cell carries only task-relevant content (intra-agent), while the scheduler selects only the best contributor per cell (inter-agent). On DAIR-V2X, the synergy is even more pronounced: the full system achieves 0.69 mAP@0.5—substantially higher than either component alone (0.55 or 0.61)—because the real-world dataset’s viewpoint diversity makes both forms of redundancy elimination jointly beneficial.
Effect of Redundancy. Theorem 3.1 predicts that selecting more than the Top-1 agent per cell is suboptimal under our utility proxy. Fig. 8 validates this empirically: a Top-2 policy that transmits features from the two highest-utility agents per cell is worse than Top-1 at nearly every bandwidth point on both datasets. This “more is worse” effect stems from a limitation of current fusion algorithms dominantly adopted by SOTA CP algorithms (e.g. max-pooling, attention [xu2022opv2v, hu2022where2comm, f_cooper]). The fused feature can be expressed as a weighted sum of the features from different agents. If the second-best agent’s feature is noisier or semantically inconsistent, it dilutes the clean, high-utility representation rather than reinforcing it. This corroborates JigsawComm’s design of selecting only the highest-quality, non-redundant features.
@@ 5 Discussion
Practical V2X Deployment. JigsawComm maps naturally onto the established dual-channel V2X architecture [FCC_DSRC, CV2X_ETSI_EN_303_613]: meta utility maps on the control channel, sparse BEV features on the data channel via the Collective Perception Message [cpm]. The scheduler’s utility-ranked priorities integrate directly into both scheduling-based (e.g. C-V2X [CV2X_ETSI_EN_303_613]) and contention-based protocols (e.g. DSRC [FCC_DSRC]), supporting centralized (V2I) and decentralized (V2V) modes for participants in the same communication range.
Limitations. The optimality transmission guarantees (Theorems 3.13.2) hold under the learned utility proxy as a data-driven intermediate representation, not the true task-level objective (which is unavailable during inference for optimizing the transmission decision); their practical value thus depends on how well the proxy tracks perception accuracy (validated empirically in Sec. 4). The current design assumes a shared BEV backbone across agents; heterogeneous architectures would require feature alignment before utility comparison [lu2024heal]. Finally, the “more is worse” finding (Sec. 4.4) highlights a limitation of existing fusion modules [f_cooper, xu2022opv2v, xu2022v2x, hu2022where2comm]; developing fusion architectures that can leverage multi-agent redundancy synergistically remains an open problem.
@@ 6 Conclusion and Future Work
[…]
Future Work. Several promising directions remain. First, the sparse features produced by JigsawComm are amenable to entropy coding, which could further reduce the transmitted payload. Second, extending the framework to joint source-channel coding would enable end-to-end optimization over unreliable wireless channels rather than treating compression and transmission separately. Third, adapting the framework to camera-based BEV features and multi-modal (LiDAR + camera) fusion would broaden its applicability to the heterogeneous sensor suites.
Supplementary Material
Appendix 0.A Proof to Theorem 3.2
Proof
By Theorem 3.1, each cell yields at most one admissible candidate; thus the global problem is $\max_{\mathcal{S}\subseteq\mathcal{C}}\ \sum_{(i,l)\in\mathcal{S}}u_{i}^{l}\;\text{s.t. }\ |\mathcal{S}|\leq K,$ where $\mathcal{C}$ is the set of at-most-one-per-cell candidates and all costs equal $c$. This is a cardinality-constrained selection problem. Let $\mathcal{G}$ be the greedy solution: the $K$ items with the largest utilities. Assume for contradiction that $\mathcal{G}$ is not optimal. Let $\mathcal{O}$ be an optimal solution with $|\mathcal{O}|\leq K$ and $\sum_{o\in\mathcal{O}}u(o)>\sum_{g\in\mathcal{G}}u(g)$. If $|\mathcal{O}|<K$, augment $\mathcal{O}$ by adding the highest-utility items not in $\mathcal{O}$ until its size reaches $K$; this cannot decrease its objective value because utilities are nonnegative, so w.l.o.g. take $|\mathcal{O}|=K$. Order items so that $u(g_{1})\geq\cdots\geq u(g_{K})$ and $u(o_{1})\geq\cdots\geq u(o_{K})$. Since $\mathcal{G}$ contains the top-$K$ utilities overall, $u(g_{j})\geq u(o_{j})$ for every $j=1,\dots,K$. Summing yields $\sum_{j}u(g_{j})\geq\sum_{j}u(o_{j})$, a contradiction. Hence $\mathcal{G}$ is optimal.
Table 3: Detailed architectural specifications for JigsawComm. NN is the number of cooperative agents in a scene, $P$ the number of non-empty pillars, and $A$ the number of anchors per cell. With voxel resolution $0.4$ m and the default LiDAR range, the initial BEV grid is $H_{0}{\times}W_{0}{=}200{\times}704$; after the multi-scale backbone the feature map is $H{\times}W{=}100{\times}352$.
Output size JigsawComm framework
[…]
Pillar Scatter $N\times 64\times H_{0}\times W_{0}$ Scatter pillar features to BEV grid
[…]
BEV Backbone (decoder) $N\times 384\times H\times W$
[…]
Concat3, 384
Sparse Thresholding $N\times 384\times H\times W$ $\mathbf{F}_{i}\leftarrow\mathbf{F}_{i}\odot\mathbb{I}[\mathbf{F}_{i}>\kappa]$ ,   $\kappa$ learnable
[…]
Transmission Scheduler $N\times 1\times H\times W$
Train:  $\tilde{m}_{i}^{l}=\sigma\!\big(\tfrac{u_{i}^{l}-\tau}{\eta}\big)\cdot\boldsymbol{\beta}_{i}^{l}(\gamma)$(STE)
Infer:  $m_{i}^{l}=\mathbb{I}[i{=}\arg\max_{k}u_{k}^{l}]\cdot\mathbb{I}[u_{i}^{l}{\geq}\tau]$
Budget:  greedy top-$K$ by $u_{i}^{l}$
Fusion $1\times 384\times H\times W$ Element-wise max over $N$ agents
Detection Head $H\times W\times A$
[…]
Appendix 0.B Architectural Configurations
Given the definitions in Sec. 3, the entire JigsawComm pipeline can be concisely formulated as follows:
$\mathbf{z}_{i}$ $=\mathrm{PillarVFE}(X_{i}),$ $\mathbf{z}_{i}$ $\in\mathbb{R}^{P_{i}\times 64},$ $\text{for agent }i$ (12)
$\mathbf{F}_{i}$ $=\Phi_{\text{enc}}\!\big(\mathrm{Scatter}(\mathbf{z}_{i})\big),$ $\mathbf{F}_{i}$ $\in\mathbb{R}^{C\times H\times W},$ (13)
$\mathbf{F}_{i}$ $\leftarrow\mathbf{F}_{i}\odot\mathbb{I}[\mathbf{F}_{i}>\kappa],$ (14)
$u_{i}^{l}$ $=\mathrm{ReLU}\!\big(\mathbf{w}^{\top}\mathbf{f}_{i}^{l}+b\big),$ $\mathbf{U}_{i}$ $\in\mathbb{R}^{1\times H\times W},$ (15)
$m_{i}^{l}$ $=\mathrm{Sched}\!\big(\{u_{j}^{l}\}_{j=1}^{N},\,\tau,\,B\big),$ $\mathbf{M}_{i}$ $\in\{0,1\}^{H\times W},$ (16)
$\hat{Y}$ $=\Phi_{\text{dec}}\!\Big(\Phi_{\text{fuse}}\!\big(\{\mathbf{M}_{i}\odot\mathbf{F}_{i}\}_{i=1}^{N}\big)\Big),$ (17)
where $X_{i}$ denotes the raw LiDAR point cloud captured by agent $a_{i}$. These points are voxelized and fed into the Pillar VFE [lang2019pointpillars], yielding $P_{i}$ pillar features $\mathbf{z}_{i}$ of dimension $64$. The pillars are scattered onto a 2D BEV grid of size $H_{0}\times W_{0}$ and processed by the multi-scale BEV backbone encoder $\Phi_{\text{enc}}$, which consists of three downsampling blocks with $3$, 55, and $8$ convolutional layers at $64$, $128$, and $256$ channels respectively (each with stride-2 entry, BatchNorm, and ReLU), followed by three transposed-convolution upsampling blocks whose outputs are concatenated to produce C=384C{=}384-channel features $\mathbf{F}_{i}$ at half the input spatial resolution ($H{=}H_{0}/2$, $W{=}W_{0}/2$).
The sparse thresholding (Eq. 14) zeros out activations below the learnable threshold $\kappa$, producing compact features with exact structural sparsity. The FUE head (Eq. 15) is a single $1{\times}1$ convolution that maps each $C$-dimensional feature cell to a scalar utility $u_{i}^{l}$, forming the meta utility map $\mathbf{U}_{i}$. After exchanging the sparse utility maps among agents, the transmission scheduler (Eq. 16) deterministically computes the binary mask $\mathbf{M}_{i}$ via the per-cell top-1 selection, threshold $\tau$, and budget constraint $B$ (Sec. 3.4). The masked features $\mathbf{M}_{i}\odot\mathbf{F}_{i}$ are fused via element-wise max pooling across agents, and the fused representation is passed through two $1{\times}1$ convolution heads for classification ($A$ anchors) and bounding-box regression ($7A$ outputs).
The detailed layer-by-layer specifications are given in Table 3. Notably, JigsawComm introduces two lightweight additions beyond the standard PointPillar backbone [lang2019pointpillars]: (i) the FUE head (C+1=385C{+}1=385 parameters) and (ii) two learnable scalar thresholds $\kappa$ and $\tau$. The transmission scheduler is purely algorithmic with no learnable parameters. The additional parameter overhead is therefore negligible ($0.01\%$ of total model parameters).
Appendix 0.C Training Procedure
Optimization. We use Adam [kingma2014adam] with initial learning rate $2{\times}10^{-3}$, weight decay $10^{-4}$, and $\epsilon{=}10^{-10}$. The learning rate is decayed by a factor of $0.1$ at epochs $15$ and $30$ using a multi-step schedule. All models are trained for $50$ epochs with batch size $4$ on a single NVIDIA RTX 4090 GPU. We save checkpoints every $2$ epochs and select the model with the lowest validation loss. We select $\lambda=1000$ for OPV2V and $\lambda=100$ for DAIR-V2X, through standard cross validation.
Temperature Annealing. The Gumbel–Softmax temperature $\gamma$ and importance-gate temperature $\eta$ share the same annealing schedule: both are initialized to $0.9$ and multiplied by a decay factor of $0.9$ each epoch, yielding $\eta_{t}=0.9^{t+1}$ at epoch $t$. The temperature is clamped at a minimum of $0.01$ to maintain numerical stability. As $\eta\to 0$, the differentiable masks converge to the deterministic inference-time policy (Prop. 1).
Data Augmentation. During training, we apply three standard augmentations to the global point cloud: (i) random flip along the $x$-axis, (ii) random rotation uniformly sampled from $\pi/4,\pi/4]$, and (iii) random scaling with a factor uniformly sampled from $[0.95,1.05]$.
Bandwidth Control. During training, no explicit bandwidth constraint is imposed ($B{=}\infty$); the semantic loss $\mathcal{L}_{\text{semantic}}$ and the learnable threshold $\tau$ naturally drive the model toward compact transmission. At inference, the budget $B$ can be set to any desired level: the scheduler simply admits cells by descending utility until the budget is met (Sec. 3.4). This allows a single trained model to operate at any point on the bandwidth–accuracy trade-off curve without retraining.
$p_{\text{drop}}$ AP@.3 AP@.5 AP@.7
0.00 0.933 0.925 0.826
0.05 0.930 0.922 0.822
0.10 0.926 0.919 0.817
0.15 0.923 0.915 0.813
0.20 0.919 0.911 0.808
Table 4: Robustness to packet loss on OPV2V. $p_{\text{drop}}$: per-agent drop probability; ego fallback at affected cells.
Appendix 0.D Scaling with More Agents
To further test JigsawComm’s ability to scale beyond the 2-5 agents available from the OPV2V and DAIR-V2X dataset Sec. 4.1, we conducted further experiments on a custom collected dataset. We utilized the CARLA digital driving simulator [Dosovitskiy17Carla] in conjunction with the OpenCDA framework [opencda] for coordinated multi-agent simulation and data recording. The data was collected following the same format and conventions as the public OPV2V benchmark dataset, including sensor configurations, coordinate systems, and CAV parameters, to ensure compatibility with existing CP models and evaluation pipelines. We set up the scene within CARLA’s ‘Town10HD’ map with 10–20 CAVs following realistic traffic rules, at a large multi-lane intersection.
Fig. 9 presents the result of JigsawComm on the custom dataset. We can observe that JigsawComm’s total data-channel payload does not double when scaling from 10 to 20 agents. It saturates and remains relatively flat when the number of agents is larger than 15, confirming the asymptotic $\mathcal{O}(1)$ data-channel scaling. As more agents join, the collective coverage saturates, the overlap becomes dominant, and the marginal contribution of each new agent becomes smaller.
Appendix 0.E Robustness to Packet Loss
In real-world V2X deployments, transmitted feature payloads may be lost due to channel fading, congestion, or collisions at the MAC layer. We study how JigsawComm behaves when a non-ego agent’s scheduled feature packet is dropped with probability $p_{\text{drop}}\in\{0,0.05,0.10,0.15,0.20\}$.
Fallback mechanism. Under the top-1-per-cell policy, each spatial cell is assigned to exactly one agent. When a non-ego agent’s packet is lost, the ego vehicle falls back to its own local feature at the affected cell. Since the ego’s feature is always available (no transmission required), this fallback incurs zero additional communication and requires no retraining. Alternatively, the system can opt for Top-$K$ scheduling when the bandwidth is sufficient.
Results. Table 4 reports the results on OPV2V. JigsawComm exhibits remarkable resilience: even at a 20% packet drop rate, AP@0.5 decreases by only 1.4 percentage points (from 0.925 to 0.911), and AP@0.7 by 1.8 points (from 0.826 to 0.808). The degradation is nearly linear in $p_{\text{drop}}$, confirming that there is no catastrophic failure mode. This resilience stems from two properties of JigsawComm’s design: (i) the top-1 scheduler assigns each cell to the single highest-utility agent, so the dropped feature is typically from a non-ego agent that only marginally outperforms the ego’s own observation at that cell (due to less occlusion or high-quality observation); and (ii) the ego’s local feature, while not the top-ranked choice, still provides a semantically meaningful representation for non-occluded objects.
Top-$K$ as an alternative. An alternative strategy for handling packet loss, especially for enhancing the detection of occluded objects, is to relax the top-1 policy to a Top-$K$ policy, where the scheduler pre-selects $K$ candidate agents per cell and falls back to the next-best candidate upon packet loss. However, Top-2 scheduling can incur additional transmission overhead. Therefore, the ego-fallback mechanism is preferable: it achieves the full accuracy benefit of Top-1 scheduling when packets arrive, and gracefully falls back to the ego’s own representation otherwise—all without increasing the nominal bandwidth or requiring changes to the scheduling policy.
Appendix 0.F Qualitative Results
JigsawComm ERMVP Where2comm
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 10: Qualitative comparison of 3D detection results on OPV2V. Each row shows one scene; columns correspond to JigsawComm, ERMVP [Zhang_2024_CVPR], and Where2comm [hu2022where2comm]. Green: ground truth; Red: predictions. JigsawComm produces tighter bounding boxes with fewer missed detections, especially for distant or partially occluded vehicles.